OpenStudy (anonymous):

My answer is C Solve 7.74x + 3 = 10.2 (2 points) Answer x ≈ -2.15 x ≈ -0.47 x ≈ 0.03 x ≈ 0.28

OpenStudy (anonymous):

are your decimal points in the right spot?

OpenStudy (anonymous):

\[ 7.74x + 3 = 10.2\] \[x = 0.930233\]

thomaster (thomaster):

7.74x + 3 = 10.2 7.74x = 7.2 7.2 / 7.74 = 0.93

OpenStudy (anonymous):

7.74x + 3 = 10.2 7.74x = 7.2 x = 7.2/7.74=.93 I got that answer too

thomaster (thomaster):

you probably typed the answer at C wrong :P